June 14, 1999, evening shift summary

Peter Shanahan (shanahan@fnal.gov)
Tue, 15 Jun 1999 00:30:25 -0500

Shift Summary for Eve shift, 14/Jun/99

On Shift: PSH, SET

Hours of beam delivered at all intensities: 3
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0.5
Hours of special runs (laser scans, muon runs...): 0.5

Accounting of lost time: Tevatron problems, a few magnet ps problems.

========================================================================
Runs ended this shift:
Run 13593 - NA - COSMICS
Run 13594 - NA - COSMICS
Run 13595 - NA - BEAM832 (very little beam)
Run 13598 - 371 K - MUON
Run 13601 - - KTEV_PED

Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift:
Ron Ray/ BobT/ PSH - 2 accesses to replace MF2 controler.
Elizabeth/ Rick K - Replace 4413 in Sta. 3.

========================================================================
Overview of shift:
Accesses and sporadic beam. Beam intensity still too low for efficient
left/right tuning by Rick Ford.

Smell of overheated insulation traced to Crate 16 in RR 191.
The smell abated after a few minutes, but -12 V is reading out near 0. We
leave it on for now, since -12 V isn't used by any modules therein, but
this should be resolved ASAP (tomorrow day shift.) This may mean
swapping the crate and/or power supply.

Vertical trim NM2V PS was off for several hours, causing vetical
target position to be off by ~1mm. We took no E832 data in this
condition.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13592 99999 1600 RSK remade l3_drivers with update reg mon
13593 1615 MF2 dropped out of laser illumination.
13593 1700 RSK/SET swap 4413 for RC10 (Sta. 3)
13594 1800 PSH/BobT swap 2 more 1445 HV controlers for MF2
13595 1930 "Hot insulation" smell detected. Traced to RR191
2000 RSK remade 2pi0 yield monitoring
2030 RSK unplugged unused DC sources to reduce data volume
2115 Hogan installs 50usec dummy busy to study DA induced
DC noise. This is still in at end of shift.
2125 NM4AN set script died at 1400A. Manually set to 1600
2210 SET tweaks ~20 Reg HVs.
13601 99999 2305 Fera peds updated following this ped run.
13601 99999 2310 NM2V PS fixed. Beam has correct vertical position.
13602 1 2312 Start Beam832 run with low intensity
13602 7 2323 Quench at B4. No beam 20 minutes

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Please watch Camac crate 16 in RR 191 closely. CAMAC crate and/or
power supply may need to be swapped during day shift.